Parth (parthkohli):

Practising for the substitution rule. The question is as follows:\[\int {x \over \sqrt{1 - 4x^2}}dx \]I need you guys to check each of my step one-by-one. Thank you.

Parth (parthkohli):

Okay, so first I take \(u = \sqrt{1 - 4x^2}\)

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

it's better to take u = 1 - 4x^2 actually

Parth (parthkohli):

Then \(du = -8xdx\)

OpenStudy (anonymous):

no..wait...remove square root

Parth (parthkohli):

Sorry...

OpenStudy (lgbasallote):

proceed

Parth (parthkohli):

Ok, then\[\int x(1 - 4x^2)^{-1 \over 2} dx \]

OpenStudy (lgbasallote):

mmhmm

Parth (parthkohli):

Oh, and \(xdx = {-1 \over 8}du\)

Parth (parthkohli):

Right?

OpenStudy (lgbasallote):

yes...

Parth (parthkohli):

\[\int u^{-1 \over 2} xdx \]

OpenStudy (lgbasallote):

mmhmm

OpenStudy (chihiroasleaf):

substitute the xdx

Parth (parthkohli):

\[\int u^{-1 \over 2} {-1 \over 8}du \]

OpenStudy (lgbasallote):

go on...

Parth (parthkohli):

\[{-1 \over 8}\int u^{-1 \over 2} du \]

OpenStudy (chihiroasleaf):

yes

OpenStudy (lgbasallote):

good

Parth (parthkohli):

\[{-1 \over 8}u^{1 \over 2} + c \]

Parth (parthkohli):

-1/4*

OpenStudy (lgbasallote):

yepp

Parth (parthkohli):

And then you substitute \(u\) back.

Parth (parthkohli):

\[{-1\over 4}(1 - 4x^2)^{1 \over 2} +c\]

OpenStudy (lgbasallote):

roght

OpenStudy (lgbasallote):

let me show you what happens if you let \(u = \sqrt{1-4x^2}\) (because im bored) \[\implies \int \frac{xdx}{\sqrt{1-4x^2}}\] let \[a = \sqrt{1-4x^2}\] \[a^2 = 1 - 4x^2\] \[2ada = -8xdx\] \[-\frac{2ada}{8} = xdx\] \[-\frac{ada}{4} = xdx\] so if you substitute... \[\implies \int (\frac 1a) (-\frac{ada}4)\] \[\implies \int -\frac{da}4\] \[\implies -\frac a4 + C\] sub back... \[\implies -\frac{\sqrt{1 - 4x^2}}4 + c\] same shiz

OpenStudy (lgbasallote):

this is actually a method called algebraic substitution (simpler variation) that is very similar to u-sub in this scenario..bu usually they are very different

Parth (parthkohli):

lol, same shiz.

OpenStudy (lgbasallote):

it's the only expression i have left that hasn't been commercialized...now that <tips hat> is gone